# MAPS: A Multi-Agent Framework Based on Big Seven Personality and Socratic Guidance for Multimodal Scientific Problem Solving

### v1

@@ Appendix A Datasets and Tasks
This study utilizes the latest three MSPs datasets, namely MathVista [23], OlympiadBench [16], and EMMA [14].
Tasks ABBR. Test
@@
Appendix B Experiment Settings and Baselines
We select GPT-4o [^3] https://openai.com/api/, a powerful MLLM, as our primary agent for solving MSPs tasks. GPT-4o not only demonstrates strong reasoning and generation capabilities across a wide range of multimodal processing tasks, but also excels in efficiently exploring multiple perspectives when faced with complex scientific domain requirements. This makes it well-suited for adaptation to various tasks and datasets within the MAPS process. We use accuracy as our primary evaluation metric to comprehensively assess the performance of different methods across diverse task scenarios. The experimental results present in Table 1 offer a thorough comparison of the performance of MAPS against all baseline methods. The experiments are primarily conducted on three MSPs datasets in  Appendix A, where we provide a detailed comparison of MAPS and the baseline models across four types of tasks: mathematics, physics, chemistry, and general tasks. Our approach achieves a new SOTA performance. To further strengthen the comparison,  Appendix C includes a generalization experiment conducted on the physics data subset of the DiagramQG [^4] https://huggingface.co/datasets/zhibei1204/DiagramQG, which further demonstrates the robustness and effectiveness of our model.
@@ Appendix B Experiment Settings and Baselines
We compare MAPS with three categories of baseline methods: original baselines, direct approach-based strong baselines, and CoT-enhanced strong baselines. Specifically, (1) the original category refers to two methods: random selection and human expert selection. These methods provide two distinct original baselines—one based on randomness and the other based on authority. (2) The direct approach-based strong baselines include some of the most powerful closed-source and open-source large language models (MLLMs) currently available worldwide. These include Claude 3.5 Sonnet [^5] https://www.anthropic.com/news/claude-3-5-sonnet, Gemini 2.0 Flash [^6] https://deepmind.google/technologies/gemini/flash/, GPT-4o, Qwen2.5-VL-72B [^7] https://huggingface.co/Qwen/Qwen2.5-VL-72B-Instruct, InternVL2.5-8B-MPO [^8] https://huggingface.co/OpenGVLab/InternVL2-8B-MPO, and LLaVA-Onevision-72B [^9] https://huggingface.co/llava-hf/llava-onevision-qwen2-72b-ov-hf. (3) To ensure a fair comparison, the third category of baselines builds on the second by adding Chain-of-Thought (CoT) reasoning, which aims to enhance the capabilities of the strong MLLMs from the second category [38].
Appendix C Generalization Experiments
@@
Across various question categories in the DiagramQG physical dataset, MAPS achieved a significant performance boost, with the maximum improvement reaching 19.51% and an overall improvement of 7.71%. These results not only demonstrate the superiority of MAPS but also indicate that its Seven Personality Agents architecture has a strong generalization ability, enabling it to maintain excellent performance across different datasets and tasks. This provides strong support for the practical application of MAPS, showcasing its potential in tackling complex tasks.
Appendix D Full-Process of MAPS
@@
Figure 9: A complete example of the collaborative output from all agents in an iteration, using the multimodal physics problem from the OlympiadBench dataset. This example demonstrates how each agent contributes to the problem-solving process, collaborating to produce a refined solution step by step.
Figure 9 presents a comprehensive step-by-step process for solving a MSPs task, using a multimodal physics problem from the EMMA dataset. As illustrated in  Figure 3, this example demonstrates the entire process of solving a physics problem. By utilizing MAPS agents based on the Big Seven Personality Model, the system engages in collaborative learning, progressively solving the problem and arriving at the correct final answer.
From the Manager, which specifically plans the entire MSPs solving process and feedback strategies, to the UserProxy that accepts inputs and task descriptions, the solution process is carried out in four key steps. These include the Interpreter, Aligner, Scholar, and Solver, each responsible for gradually refining the solution step-by-step. The Critic agent ultimately evaluates and provides feedback on these four steps, pinpointing the areas requiring modification and backtracking to correct the solution process.  Figure 9 sequentially illustrates the roles of each agent and their collaboration, showcasing how problem solving and optimization are effectively executed at every stage, ensuring both the accuracy and efficiency of the final result.
Appendix E Prompts for Agents
Tables 4, 5 and 6 provides a summary of the prompts used for each agent in this paper, with each agent playing a pivotal role in the overall problem-solving process. The process begins with the Manager, which is responsible for strategically planning the entire MSPs solving procedure and feedback mechanisms. The Manager defines the high-level flow, ensuring that all agents operate cohesively. The UserProxy agent then takes over by accepting the user inputs and task descriptions, forming the basis for the entire problem-solving process.
Following this, the solution is gradually refined by four core agents: Interpreter, Aligner, Scholar, and Solver. The Interpreter first processes the task description, breaking down and understanding the problem. The Aligner ensures that the problem is mapped to the correct framework and available tools. The Scholar conducts any necessary research and gathers knowledge from relevant sources, while the Solver works through the problem systematically, progressively generating solutions. Once the initial solution is formed, the Critic agent assesses each of the four previous steps, providing feedback on areas that need refinement. The Critic then suggests modifications and backtracks to revise the solution, ensuring the process remains optimized and the final result is both accurate and robust.
Manager
You are a task manager, responsible for managing and deciding the order of each Agent’s output, ensuring that tasks are executed in the correct sequence. You will determine when each step should be executed based on the task requirements and coordinate the outputs of each Agent. Your goal is to ensure that the task process is efficient and orderly, and adjust the execution priority of steps when necessary.
UserProxy
Here are some questions that require careful thought, so please think deeply, solve the questions carefully, and output your answers.
@@ Appendix E Prompts for Agents
Table 4: A summary of the prompts used by the Manager, UserProxy, Interpreter, and Aligner agents in this paper.

### v2

@@ Appendix B Datasets
This study utilizes the latest three multimodal reasoning datasets, namely MathVista, OlympiadBench, and EMMA.
Dataset ABBR. Size
@@
Appendix C Experiment Settings and Baselines
We select GPT-4o (achiam2023gpt), a powerful MLLM, as our primary agent for complex problem reasoning. GPT-4o not only demonstrates strong reasoning and generation capabilities across a wide range of multimodal processing tasks, but also excels in efficiently exploring multiple perspectives when faced with complex scientific domain requirements. This makes it well-suited for adaptation to various tasks and datasets within the MAPS process. We use accuracy as our primary evaluation metric to comprehensively assess the performance of different methods across diverse task scenarios. The experimental results offer a thorough comparison of the performance of MAPS against all baseline methods. The experiments are primarily conducted on three datasets in Appendix B, where we provide a detailed comparison of MAPS and the baseline models across four types of tasks: mathematics, physics, chemistry, and general tasks. Our approach achieves a new SOTA performance. To further strengthen the comparison, Appendix D includes a generalization experiment conducted on the physics data subset of the DiagramQG (zhang2025diagram), which further demonstrates the robustness and effectiveness of our model.
We compare MAPS with three categories of baseline methods: original baselines, direct approach-based strong baselines, and CoT-enhanced strong baselines. Specifically, (1) the original category refers to two methods: random selection and human expert selection. These methods provide two distinct original baselines—one based on randomness and the other based on authority. (2) The direct approach-based strong baselines include some of the most powerful closed-source and open-source large language models (MLLMs) currently available worldwide. These include Claude 3.5 Sonnet (kevian2024capabilities), Gemini 2.0 Flash (comanici2025gemini), GPT-4o, Qwen2.5-VL-72B (bai2023qwen), InternVL2.5-8B-MPO (chen2024expanding), and LLaVA-Onevision-72B (li2024llava). (3) To ensure a fair comparison, the third category of baselines builds on the second by adding Chain-of-Thought (CoT) reasoning, which aims to enhance the capabilities of the strong MLLMs from the second category.
Appendix D Generalization Experiments
@@
Across various question categories in the DiagramQG physical dataset, MAPS achieved a performance boost, with the maximum improvement reaching 19.51% and an overall improvement of 7.71%. These results not only demonstrate the superiority of MAPS but also indicate that its Five Personality Agents architecture has a strong generalization ability, enabling it to maintain excellent performance across different datasets and tasks. This provides strong support for the practical application of MAPS, showcasing its potential in tackling complex tasks.
Appendix E Full Process of MAPS
@@
Figure 10: A complete example of collaborative output from all agents in an iteration, using a multimodal physics problem from OlympiadBench involving the axial electric field of a dipole in the far-field. This illustrates how agents work together step by step to refine the solution.
Figure 10 presents a comprehensive step-by-step process for complex problem reasoning, using a multimodal physics problem from the EMMA dataset. This example demonstrates the entire process of solving a physics problem. By utilizing MAPS agents based on the Big Five Personality Model, the system engages in collaborative learning, progressively solving the problem and arriving at the correct final answer.
The solution process is carried out in four key steps. These include the Interpreter, Aligner, Scholar, and Solver, each responsible for gradually refining the solution step-by-step. The Critic agent ultimately evaluates and provides feedback on these four steps, pinpointing the areas requiring modification and backtracking to correct the solution process. Figure 10 sequentially illustrates the roles of each agent and their collaboration, showcasing how problem solving and optimization are effectively executed at every stage, ensuring both the accuracy and efficiency of the final result.
Appendix F Prompts for Agents
Table 45, and 6 provides a summary of the prompts used for each agent in this paper, with each agent playing a pivotal role in the overall problem-solving process.
The solution is gradually refined by four core agents: Interpreter, Aligner, Scholar, and Solver. The Interpreter first processes the task description, breaking down and understanding the problem. The Aligner ensures that the problem is mapped to the correct framework and available tools. The Scholar conducts any necessary research and gathers knowledge from relevant sources, while the Solver works through the problem systematically, progressively generating solutions. Once the initial solution is formed, the Critic agent assesses each of the four previous steps, providing feedback on areas that need refinement. The Critic then suggests modifications and backtracks to revise the solution, ensuring the process remains optimized and the final result is both accurate and robust.
@@ Appendix F Prompts for Agents
Table 4: A summary of the prompts used by the Interpreter, and Aligner agents in this paper.